A 2.4 kg pan at 155 degrees C is dumped into 3.5 kg of water at 15 degrees C. What is the change in entropy? Specific heat of pan is 900 J/(kg*K), and specific heat of water is 4184 J/(kg*K).
Answer #1

albot - From heat balance mp Cp (tp-t) = mau Cw (t-tw) 2,4*9oo * (155-t) =3.5*4184 (t-15) *2160 (155-t) = 14644 (t-15) » 3348

306 - 5°C = 288k T 288 for Water Asw Smw ew di = 3.5*4184 dat T] = 3:5*4184 [ln 3o6-ln 288] =887.78 JIK 306 288 Thas change i